OpenStudy (gudetamas):

Write a function that has a vertex of (2,4) and contains the point (4,16)

OpenStudy (518nad):

hello so take a parabola

OpenStudy (gudetamas):

i got y=3(x+2)^2+4

OpenStudy (518nad):

your basic parabola with slope m has form y=mx^2 if u want to move it to ur vertex y-4 = m(x-2)^2 now if u will pick an m such that your point is satisfied

OpenStudy (518nad):

its x-2 not x+2, x+2 will make it shift to -2

OpenStudy (gudetamas):

ok, is everything else good?

OpenStudy (518nad):

k find slope now

OpenStudy (518nad):

y-4 = m(x-2)^2 point (4,16) exists on parabola 16-4=m(16-2)^2 solve for m

OpenStudy (gudetamas):

3?